# TPK: Trustworthy Trajectory Prediction Integrating Prior Knowledge For Interpretability and Kinematic Feasibility

### v2

@@ V-C1 Ablation Study
First, we assess the impact of our modifications on the network’s accuracy. Table IV further confirms our decision for GnL over MnR and DG-SFM over SKGACN, as the networks using these methods also lead to a higher accuracy. Moreover, the step-wise integration of our modifications shows that each layer enhances the model performance after an initial deterioration due to the kinematic layer, which reflects the accuracy limitations arising from the dataset’s infeasible ground truth.
@@ V-C2 Comparison to the State-of-the-Art
Finally, we compare our final model with the official HPTR model. As shown in Table V, the official HPTR model shows better accuracy than our model. However, this comparison is distorted by differences in the training loss function and epoch counts. The official HPTR uses a negative log-likelihood loss, which has an edge over our Huber loss in learning multimodal distributions through Gaussian position representation [35]. However, adapting this loss for feasibility-constrained networks lies beyond the scope of this work. Moreover, the official HPTR’s 75-epoch training period, compared to the 30 epochs of our model, gives it a notable advantage. Extending our training to 75 epochs was not viable due to the 30-day time requirement, which exceeded our resource constraints.
Our ablation results for the different prior integration methods in Table IV closely matched plain HPTR, suggesting that the remaining performance gap is almost exclusively due to the kinematic layer. However, this trade-off is well justified. As discussed in Section V-B1, the kinematic layer prevents the model from learning unrealistic and noisy patterns present in the dataset. On balance, our model outweighs its minor accuracy disadvantage with two crucial benefits: more trustworthy reasoning through interpretable interaction attention, and more trustworthy prediction through guaranteed physical feasibility.

### v3

@@ V-C1 Ablation Study
First, we assess the impact of our modifications on the network’s accuracy. Table˜IV further confirms our decision for GnL over MnR and DG-SFM over SKGACN, as the networks using these methods also lead to a higher accuracy. Moreover, the step-wise integration of our modifications shows that each layer enhances the model performance after an initial deterioration due to the kinematic layer, which reflects the accuracy limitations arising from the dataset’s infeasible ground truth.
@@ V-C2 Comparison to the State-of-the-Art
Finally, we compare our final model with the official HPTR model. As shown in Table˜V, the official HPTR model shows better accuracy than our model. However, this comparison is distorted by differences in the training loss function and epoch counts. The official HPTR uses a negative log-likelihood loss, which has an edge over our Huber loss in learning multimodal distributions through Gaussian position representation [35]. However, adapting this loss for feasibility-constrained networks lies beyond the scope of this work. Moreover, the official HPTR’s 75-epoch training period, compared to the 30 epochs of our model, gives it a notable advantage. Extending our training to 75 epochs was not viable due to the 30-day time requirement, which exceeded our resource constraints.
Our ablation results for the different prior integration methods in Table˜IV closely matched plain HPTR, suggesting that the remaining performance gap is almost exclusively due to the kinematic layer. However, this trade-off is well justified. As discussed in Section˜V-B1, the kinematic layer prevents the model from learning unrealistic and noisy patterns present in the dataset. On balance, our model outweighs its minor accuracy disadvantage with two crucial benefits: more trustworthy reasoning through interpretable interaction attention, and more trustworthy prediction through guaranteed physical feasibility.